OpenStudy (anonymous):

Solve the Quadratic Formula. x^2+5x+6=0 Btw the ^2 is x with a two on top for squared. x=__and__

OpenStudy (anonymous):

Use the quadratic formula do you know what that is?

OpenStudy (anonymous):

Or you can complete the square.

OpenStudy (anonymous):

No thats why i asked(:

OpenStudy (anonymous):

I dont knw

OpenStudy (anonymous):

\[\begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & {{\rm{when}}} & {ax^2 + bx + c = 0} \\ \end{array} \]

OpenStudy (anonymous):

There use that!!

OpenStudy (anonymous):

I dont know )PLZ just help me

OpenStudy (anonymous):

WHat are u confused about?

OpenStudy (anonymous):

I just dont understand it?

OpenStudy (anonymous):

look at the right equation that i posted It looks like the one u posted right? Find out what a b and c are equal to and plug them in to the left equation and solve for x

OpenStudy (anonymous):

I can't helpp u if you dont tell me what part you are confused about

OpenStudy (anonymous):

Can y0uo just tell me the answer this one time? Its summer and im trying tohelp my little brother, he wants to be smart for junior high aha

OpenStudy (anonymous):

Hello??